OpenStudy (anonymous):

Which of these has the greatest area? Use 3.14 for pi. A) rhombus with side lengths: 9 inches, height: 4 inches B) triangle with base length: 8 inches, height: 3 inches C) circle with radius: 2 inches D) trapezoid with base lengths: 5 inches and 3 inches, height: 4 inches

OpenStudy (anonymous):

A) A = bh = 9(4) = 36 B) A = 1/2bh = 1/2(8)(3) = 12 C) A = πr^2 = (3.14)(4) = 12.56 D) A = h/2(b_1 + b_2) = 2(5 + 3) = 2(8) = 16 Which one is biggest?

OpenStudy (anonymous):

oh thanks!

OpenStudy (anonymous):

You're welcome :)